### v1

@@ Acknowledgements.
The authors thank the organizers of the Pollica 2023 SIDM Workshop, where part of this work was done. In addition, the authors thank all participants of the Darkium SIDM Journal Club for discussions. MSF and LS thank Niklas Becker, Andrew Benson, Akaxia Cruz, Xiaolong Du, Frederick Groth, Charlie Mace, Annika H. G. Peter, Sabarish Venkataramani, Daneng Yang, Shengqi Yang, Hai-Bo Yu, Xingyu Zhang, and Zhichao Carton Zeng as well as the members of the CAST group at the University Observatory Munich for helpful discussions. This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094 “Origins” – 390783311. Software: NumPy (Harris et al., 2020), Matplotlib (Hunter, 2007).
@@ Appendix A Unresolved mean free path
Figure 13: The position and velocity of the BH in our test setup is shown for a simulation with a fSIDM cross-section of $\sigma_{\mathrm{V}}/m=3\times 10^{-5}\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$. The high resolution corresponds to $N=5\times 10^{6}$ particles and the low resolution to $N=1\times 10^{6}$ particles. The larger box size of $l_{\mathrm{box}}=829.5\,\mu\mathrm{pc}$ was used.

### v2

@@ Acknowledgements.
The authors thank the organisers of the Pollica 2023 SIDM Workshop, where part of this work was done. In addition, the authors thank all participants of the Darkium SIDM Journal Club for discussions. MSF and LS thank Gonzalo Alonso-Álvarez, Niklas Becker, Andrew Benson, Akaxia Cruz, Xiaolong Du, Frederick Groth, Charlie Mace, Annika H. G. Peter, Rainer Spurzem, Sabarish Venkataramani, Daneng Yang, Shengqi Yang, Hai-Bo Yu, Xingyu Zhang, and Zhichao Carton Zeng as well as the members of the CAST group at the University Observatory Munich for helpful discussions. They are also grateful to Bradley Kavanagh for helpful comments on the draft. This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094 ‘Origins’ – 390783311. Software: NumPy (Harris et al., 2020), Matplotlib (Hunter, 2007).
@@ Appendix A Unresolved mean free path
Figure 14: Position and velocity of the BH in our test setup for a simulation with a fSIDM cross-section of $\sigma_{\mathrm{V}}/m=3\times 10^{-5}\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$. The high resolution corresponds to $N=5\times 10^{6}$ particles and the low resolution to $N=1\times 10^{6}$ particles. The larger box size of $l_{\mathrm{box}}=829.5\,\mu\mathrm{pc}$ was used.